[Ncep.nhc.nco_contacts] TSB Morning Rounds - Wed Apr 19, 2017

TSB Admin nhc.tsbadmin at noaa.gov
Wed Apr 19 15:07:30 UTC 2017

Here is today's summary of NHC computer operations:
--- Craig


1. TSB Developer Matt Sardi is continuing to work with NCO Linux Sysadmins  
to fix symbolic links and disk mounts on the TAFB N-AWIPS workstations.

2. AWIPS-II Sysadmin Chris Mello repaired misaligned floor panels and  
tangled cables under the Surface Analyst's workstation.

1. ATCF issues: Problems transmitting the HWRF/GFDL 06Z CARQ, which had to  
be sent twice for the SDM to receive it. The latest submittals/retrievals  
seem to be working fine. TSB will continue to monitor the dataflow between  
the ATCF servers and WCOSS systems.

2. Sr. Hurricane Specialist Lixion Avila began issuing advisories on  
Subtropical Depression One (storm ID al012017). Some bugs were discovered  
in the TC Graphics (wwgraph ksh script) and with NHC's web site. TSB  
Developers Matt Onderlinde and Dave Zelinsky are fixing these problems.

3. Lixion also successfully submitted P-Surge test runs both yesterday and  
Other items:


-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20170419/476b3c17/attachment.html 

More information about the Ncep.nhc.nco_contacts mailing list